# Large language models are potential rote learners

### v3

@@
Large language models are potential rote learners
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
1College of Computer Science and Technology, Zhejiang University
2School of Data Science and Engineering, East China Normal University
3State Key Laboratory of Transvascular Implantation Devices,
The Second Affiliated Hospital Zhejiang University School of Medicine
4School of Public Health, Zhejiang University
5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
6Alibaba Cloud Computing
xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
Disentangling Memorization and Capability Learning in Large Language Models
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
1College of Computer Science and Technology, Zhejiang University
2School of Data Science and Engineering, East China Normal University
3State Key Laboratory of Transvascular Implantation Devices,
The Second Affiliated Hospital Zhejiang University School of Medicine
4School of Public Health, Zhejiang University
5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
6Alibaba Cloud Computing
xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
Large Language Models Are Potentially Rote Memorizer
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
1College of Computer Science and Technology, Zhejiang University
2School of Data Science and Engineering, East China Normal University
3State Key Laboratory of Transvascular Implantation Devices,
The Second Affiliated Hospital Zhejiang University School of Medicine
4School of Public Health, Zhejiang University
5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
6Alibaba Cloud Computing
xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
Large Language Models Are Both Rote Memorizer and Capacity Learner
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
1College of Computer Science and Technology, Zhejiang University
2School of Data Science and Engineering, East China Normal University
3State Key Laboratory of Transvascular Implantation Devices,
The Second Affiliated Hospital Zhejiang University School of Medicine
4School of Public Health, Zhejiang University
5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
6Alibaba Cloud Computing
xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
Disentangling Rote Memorization and Capability Learning for
Large Language Model Evaluation
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
1College of Computer Science and Technology, Zhejiang University
2School of Data Science and Engineering, East China Normal University
3State Key Laboratory of Transvascular Implantation Devices,
The Second Affiliated Hospital Zhejiang University School of Medicine
4School of Public Health, Zhejiang University
5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
6Alibaba Cloud Computing
xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
Large Language Models Could Be Rote Learners
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
1College of Computer Science and Technology, Zhejiang University
2School of Data Science and Engineering, East China Normal University
3State Key Laboratory of Transvascular Implantation Devices,
The Second Affiliated Hospital Zhejiang University School of Medicine
4School of Public Health, Zhejiang University
5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
6Alibaba Cloud Computing
xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
@@ Abstract
Multiple-choice question (MCQ) benchmarks are widely used for evaluating Large Language Models (LLMs), yet their reliability is undermined by benchmark contamination. In this study, we reframe contamination as an inherent aspect of learning and seek to disentangle genuine capability acquisition from superficial memorization in LLM evaluation. First, by analyzing model performance under different memorization conditions, we uncover a counterintuitive trend: LLMs perform worse on memorized MCQs than on non-memorized ones, indicating the coexistence of two distinct learning phenomena, i.e., rote memorization and genuine capability learning. To disentangle them, we propose TrinEval, a novel evaluation framework that reformulates MCQs into an alternative trinity format, reducing memorization while preserving knowledge assessment. Experiments validate TrinEval’s effectiveness in reformulation, and its evaluation reveals that common LLMs may memorize by rote 20.5% of knowledge points (in MMLU on average).
Large Language Models Could Be Rote Learners
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6 1College of Computer Science and Technology, Zhejiang University 2School of Data Science and Engineering, East China Normal University 3State Key Laboratory of Transvascular Implantation Devices, The Second Affiliated Hospital Zhejiang University School of Medicine 4School of Public Health, Zhejiang University 5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence 6Alibaba Cloud Computing xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
@@ 1 Introduction
Figure 1: MCQ-based LLM evaluation. We observe that LLMs tend to underperform on memorized MCQs.
@@ 4 Experiments
Q1. Is TrinEval knowledge-preserving in order to fulfill knowledge assessment?
Q2. Can TrinEval reduce memorization effects during capability evaluation?
Q3. What does TrinEval reveal about LLMs’ rote memorization and genuine capability?
@@ 4.1 Experiment Setup
Models. We utilize API-based commercial LLMs, specifically gpt-4o-2024-08-06 (GPT) (OpenAI, 2024) and qwen-max-2024-09-19 (Qwen) (Yang et al., 2024; Team, 2024) for question reformulation by TrinEval. Model evaluation is conducted on open-source LLMs due to limited budgets, and we experiment with three popular LLMs including Llama2-7B (Llama) (Touvron et al., 2023), Mistral-7B-v0.2 (Mistral) (Jiang et al., 2023), and Vicuna-v1.5-7B (Vicuna) (Zheng et al., 2023b). All the three LLMs are accessed from Huggingface and implemented with transformers library, we thus could obtain the log-probability of output token for fine-grained study. Throughout our tests, we use the default generation parameters and adopt greedy decoding to enhance reproducibility.
Benchmarks. We evaluate LLM on the widely used MMLU (Hendrycks et al., 2020) benchmark. MMLU consists of 57 subjects from areas including STEM, humanities, social sciences, and others, enabling comprehensive evaluation of LLM capacity. As there are duplicated MCQs across different subjects, we eliminate them and obtain 14,006 MCQs as the test set.
Evaluation. With commercial LLMs, we evaluate model performance by extracting answers with regular expressions. For open-source LLMs, we access the output probability of the first generated token (e.g., option IDs A/B/C/D) to obtain a quantitative performance result.
@@
We first verify whether the reformulation by TrinEval is knowledge-preserving in order to fulfill knowledge assessment. To achieve this, our primary objective is to validate that the reformulation approach (1) does not lose key information that results in previous correctly-answered questions being answered incorrectly and (2) does not introduce anomalous or unexpected information that results in inflated performance.
Figure 3: Knowledge-preserving validation for reformulation by TrinEval. 4,343 and 4,645 qualified MCQs are obtained with Qwen and GPT after reformulation. We test Qwen and GPT in these qualified subsets. The green and blue circles stand for the correctly answered MCQs in TrinEval and original formats, respectively.
(a) Evocation with Llama2 and dev-fsp.
(b) Evocation with Mistral and dev-fsp.
(c) Evocation with Vicuan and dev-fsp.
(d) Evocation with Llama2 and seq-fsp.
(e) Evocation with Mistral and seq-fsp.
(f) Evocation with Vicuan and seq-fsp.
Figure 4: The results of memorization evocation under various dataset-related information context, with blue and green curves referring to the memorization difference $\Delta F_{m}$ in the TrinEval and original formats, respectively. In the x-axis, ‘clean’, ‘meta’, ‘dev-fsp’, and ‘seq-fsp’ stand for without dataset-related context, with the name of the dataset, with few-shot prompt from the training set, and with few-shot prompt from the test set ahead of the current testing question. These curves indicate the growing memorization metric $\Delta F_{m}$ with the stronger dataset-related information in general. However, the $\Delta F_{m}$ by TrinEval under the strongest memory evocation context are consistently lower than those in the original format, e.g., ‘clean’.
Upon completing the complete TrinEval reformulation process, we ultimately obtained 4,343 MCQs and their corresponding knowledge entities, attributes, and contexts that met our criteria using Qwen, as well as 4,645 qualified MCQs and their respective triplets using GPT. We then instruct Qwen and GPT to answer these respective questions in both the original (baseline) and restated triplet form. The prompts and an MCQ example are available in Table 2 in Appendix and the results are shown in Fig. 3.
We can observe that for Qwen, 92.95% of correctly answered MCQs in the TrinEval format maintain their accuracy in the original format, while for GPT, 90.05% of qualified MCQs are answered correctly, with 95% of these maintaining accuracy in the original format. That is, for both Qwen and GPT, we can infer that the correctly answered MCQs from the qualified ones in TrinEval format can be regarded as a subset of the correctly answered MCQs with the original MCQ format. This proves that the proposed TrinEval reformulation method does not incorporate extra information that leads to additional capability of LLMs. On the other hand, the intersection MCQs between the correctly answered in two formats also make up of around 95% of the MCQs correctly answered in the original MCQ format, which proves that the TrinEval incorporates all the necessary information to answer the question. In conclusion, our TrinEval effectively retains the LLMs’ problem-solving capability compared to the original MCQ text.
@@
In this subsection, we aim to validate whether the proposed TrinEval can eliminate the unnecessary memorization of LLMs, and thus demonstrate enhanced robustness against various perturbations. To answer this question, following Bordt et al. (2024), we deliberately incorporate the dataset-related information into the context and evaluate whether the TrinEval reformulation can suppress the growing memorization level with memorization evocation of different extent and can reveal the genuine capabilities of LLMs.
We incorporate the dataset-related information into the context, i.e., the name of the dataset, and the few-shot prompt of samples within the same dataset for the memorization-evocation perturbation (see Appendix E). Here we use Llama, Mistral, and Vicuna as the tested LLMs since we access the output probabilities to compute the memorization metric $F_{m}$. As there is no specific zero point of $F_{m}$ indicating the absolute-no memorization of MCQs given an LLM, in order to better visualize the difference between the proposed TrinEval and the original MCQ baseline format, we take the $F_{m}$ with vanilla MCQ (i.e., original MCQ format without any dataset-related prompt) as the baseline and visualize the averaged difference between the $F_{m}$ of the tested format and the baseline.
(a) $F_{m}$ v.s. $F_{c}$ with Llama2 based on Qwen-extracted triplets.
(b) $F_{m}$ v.s. $F_{c}$ with Mistral based on Qwen-extracted triplets.
(c) $F_{m}$ v.s. $F_{c}$ with Vicuna based on Qwen-extracted triplets.
(d) $F_{m}$ v.s. $F_{c}$ with Llama2 based on GPT-extracted triplets.
(e) $F_{m}$ v.s. $F_{c}$ with Mistral based on GPT-extracted triplets.
(f) $F_{m}$ v.s. $F_{c}$ with Vicuna based on GPT-extracted triplets.
@@
Figure 5: The distribution of MCQs based on Memorization metric $F_{m}$ v.s. the Capability metric $F_{c}$. According to the values of $F_{m}$ and $F_{c}$, we separate the MCQs equally into five groups and visualize the distribution of MCQs with the heatmap from weak to strong.
Specifically, for the memorization-evocation permutation, we progressively enhance the prompt context for memory evocation, starting from merely providing the dataset name, to offering samples within the same dataset as few-shot prompts (including the training set of the dataset-dev, and the preceding samples adjacent to the test sample-seq), and finally to providing both as context. Here for each tested MCQ, we calculate the difference between the $F_{m}$ given the corresponding memorization evocation prompt and the $F_{m}$ with the vanilla MCQ baseline. Fig. 4 shows the curve based on the average difference of each MCQ.
From this figure, as stated by Bordt et al. (2024), we can see that $F_{m}$ is growing with the stronger dataset-related context. When providing more specific context related with the test dataset, the LLMs tend to exhibit stronger memorization of the MCQs. Specially, for all three open-source LLMs, the $\Delta F_{m}$ curve of TrinEval is below the curve of the original MCQ baseline. More importantly, the $F_{m}$ of TrinEval with the strongest memorization evocation is still below the vanilla MCQ baseline, which proves that TrinEval can effectively eliminate the memorization from LLMs.
4.4 Q3. TrinEval’s Findings on Memorization and Capability
Figure 6: Averaged distance of each MCQs between the closest $1\%$ MCQs’ embeddings. ‘Rote Memorization’ refers to MCQ within lower left $2\times 2$ squares that typically exhibits high memorization metric $F_{m}$ and low capability $F_{c}$ while the ‘Genuine Capability Learning’ stands for MCQ lies within the upper right $2\times 2$ squares with low $F_{m}$ but high $F_{c}$. Further results are shown in Appendix F
In this subsection, we aim to explicitly study the relationship between the memorization and the capability of LLMs with the metrics $F_{m}$ and $F_{c}$. As the commercial-API-based LLMs do not provide the output probability of the whole vocabulary, we mainly use the open-source LLMs to compute these two metrics. After obtaining the $F_{m}$ and $F_{c}$ of each MCQ, we separate all the qualified MCQs into 5 equal groups. Finally, we utilize the heatmap to reveal the relationship between the capability and the memorization of the tested LLMs.
As shown in Fig. 5, most of the MCQs concentrate on the lower left corner and the upper right corner of the heatmap. Specifically, for $F_{m}$ v.s. $F_{c}$ with Llama2 based on Qwen-extracted triplets, MCQs within the lower left $2\times 2$ squares and the upper right $2\times 2$ squares make up of the 38.57% of all the tested MCQs with a Pearson-correlation of -0.7755 (p-value $<0.05$), while the MCQs within the lower left and upper right $3\times 3$ squares make up of the 74.17% of all the tested MCQs with a Pearson-correlation of -0.8124 (p-value $<0.05$). For the results with Mistral based on Qwen-extracted triplets, MCQs within the lower left and upper right $2\times 2$ squares make up of the 44.90% of all the tested MCQs with a Pearson-correlation of -0.8722 (p-value $<0.05$), while the MCQs within the lower left and upper right $3\times 3$ squares make up of the 80.82% of all the tested MCQs with a Pearson-correlation of -0.8794 (p-value $<0.05$). More results are shown in Tab. 3. This evidence indicates that MCQs with lower memorization levels tend to exhibit better problem-solving capabilities of LLMs, while those with higher memorization levels are associated with reduced performance in solving tasks.
@@
Next, we hypothesize that the LLMs are potential rote learners through the human memory system, which has been characterized by two fundamental components: Long-Term Memorization (LTM) and Short-Term Memorization (STM) Shiffrin (2003). Neurobiological studies reveal that STM relies on transient synaptic protein synthesis with limited temporal persistence and functional scalability. In contrast, LTM is constructed through stabilized neuronal memory traces that constitute an enduring knowledge framework. This neural architecture not only supports STM operations as a cognitive substrate but also enables sophisticated information generalization across diverse contexts. As illustrated in Allen-Zhu and Li (2023) and Ovadia et al. (2023), LLMs trained with multiple rephrased corpus tend to perform better than LLMs trained with only the original corpus. When providing only one format of training corpus, similar to the STM system, LLMs tend to memorize the corpus at token-level rather than knowledge-level. In other words, LLMs encode these corpora at a shallow level with the original format. After questions are rephrased with methods like our proposed TrinEval, the input corpus seems connected with the known knowledge like the LTM for structured storage and enables sophisticated information generalization. We show more detailed results in Appendix C.
To further validate our hypothesis, we compute the embeddings of MCQs within the qualified MMLU dataset and average the distance between the other closest $1\%$ MCQs. We visualize the mean distance of MCQs within the lower left and upper right $2\times 2$ squares in Fig. 5. The results are shown in Fig. 6. We surprisingly find that the averaged distance of the Genuine Capability Learning MCQs (i.e., MCQs within the upper right $2\times 2$ squares) is almost half as much as the distance of the Rote Memorization MCQs (i.e., MCQs within the lower left $2\times 2$ squares). The result hints that the memorized MCQs are sparsely encoded by MCQs while the non-memorized ones share common embeddings, which is again coincident with the findings of the STM and LTM.
Though it is well believed that memorization may lead to better but cheating performance of LLMs, we prove that the more LLMs memorize, the worse they are at solving problems.

### v4

# Large Language Models Could Be Rote Learners

@@ Abstract
Multiple-choice question (MCQ) benchmarks are widely used for evaluating Large Language Models (LLMs), yet their reliability is undermined by benchmark contamination. In this study, we reframe contamination as an inherent aspect of learning and seek to disentangle genuine capability acquisition from superficial memorization in LLM evaluation. First, by analyzing model performance under different memorization conditions, we uncover a counterintuitive trend: LLMs perform worse on memorized MCQs than on non-memorized ones, indicating the coexistence of two distinct learning phenomena, i.e., rote memorization and genuine capability learning. To disentangle them, we propose TrinEval, a novel evaluation framework reformulating MCQs into an alternative trinity format, reducing memorization while preserving knowledge assessment. Experiments validate TrinEval’s effectiveness in reformulation, and its evaluation reveals that common LLMs may memorize by rote 20.5% of knowledge points (in MMLU on average).
@@ 1 Introduction
Figure 1: MCQ-based evaluation. We observe that LLMs tend to underperform on memorized MCQs.
@@ 4 Experiments
Q1. Does TrinEval effectively preserve knowledge to enable accurate knowledge assessment?
Q2. To what extent does TrinEval mitigate memorization effects during capability evaluation?
Q3. What insights does TrinEval provide into the distinction between rote memorization and genuine capability in LLMs?
@@ 4.1 Experiment Setup
Models. We employ API-accessible commercial LLMs for question reformulation in TrinEval, specifically gpt-4o-2024-08-06 (GPT) (OpenAI, 2024) and qwen-max-2024-09-19 (Qwen) (Yang et al., 2024; Team, 2024). For model evaluation, we utilize open-source LLMs to obtain token-level logits. Our experiments focus on three widely adopted models: Llama2-7B (Llama) (Touvron et al., 2023), Mistral-7B-v0.2 (Mistral) (Jiang et al., 2023), and Vicuna-v1.5-7B (Vicuna) (Zheng et al., 2023b). These models are accessed from Hugging Face and implemented with the transformers library, allowing us to compute the log-probabilities of generated tokens for fine-grained analysis. All evaluations are conducted on a single NVIDIA A100 GPU with 40GB of memory. Default generation parameters are used, and greedy decoding is applied throughout to ensure reproducibility.
Benchmarks. We evaluate LLMs on the widely used MMLU benchmark (Hendrycks et al., 2020), in light of growing evidence of data contamination (Touvron et al., 2023) [^3] https://hitz-zentroa.github.io/lm-contamination/ in many recent LLMs on this data. MMLU spans 57 subjects across STEM, humanities, social sciences, and others, providing a comprehensive assessment of model capabilities. We find duplicated MCQs across subjects on MMLU, and thus de-duplicate the dataset, resulting in a test set of 14,006 unique questions.
Evaluation. With commercial LLMs, we evaluate model performance by extracting predicted answers with regular expressions. For open-source models, we access the output probability of the first generated token (i.e., corresponding to multiple-choice option IDs A/B/C/D) to compute quantitative performance metrics.
@@
Figure 3: Knowledge-preserving validation of TrinEval reformulation. The green, blue, and overlapping regions represent the sets of MCQs correctly answered in the original format, TrinEval format, and both formats, respectively. Best viewed in color.
We begin by verify whether TrinEval’s reformulation preserves essential knowledge, a prerequisite for reliable knowledge assessment. Specifically, our goal is to ensure that (1) the reformulation does not omit critical information that would cause originally correctly answered questions to be answered incorrectly, and (2) it does not introduce anomalous or unexpected content that could artificially inflate performance.
Following the complete TrinEval reformulation pipeline, we obtain 4,343 qualified MCQs with corresponding knowledge entities, attributes, and contexts using Qwen, and 4,645 MCQs with associated triplets using GPT. We then prompt both Qwen and GPT to answer their respective MCQs in both the original and reformulated triplet-based formats. Example MCQs and the questioning prompt templates in both formats are provided in Table 2 in the Appendix; results are summarized in Fig. 3.
Among the qualified MCQs reformulated by Qwen and GPT, 4,133 and 4,409 are answered correctly in at least one of the two formats, respectively. Notably, the majority of these are answered correctly in both the original and TrinEval formats, accounting for 90.3% with Qwen and 90.4% with GPT, indicating strong consistency in terms of problem-solving across formats. Only 121 questions for Qwen (2.9%) and 226 for GPT (5.1%) are answered correctly exclusively in the original format, suggesting that the TrinEval reformulation does not omit essential information. Conversely, 278 questions for Qwen (6.7%) and 197 for GPT (4.5%) are answered correctly only in the reformulated format, which is comparable in general to the numbers for the original format alone. This further suggests that the reformulation does not introduce extra information that might artificially enhance model performance. Collectively, these findings provide strong evidence that TrinEval preserves the core knowledge necessary for accurate answering, satisfying the requirements for reliable capability evaluation.
@@
(a) Evoc. with Llama2 and dev-fsp.
(b) Evoc. with Mistral and dev-fsp.
(c) Evoc. with Vicuna and dev-fsp.
(d) Evoc. with Llama2 and seq-fsp.
(e) Evoc. with Mistral and seq-fsp.
(f) Evoc. with Vicuna and seq-fsp.
Figure 4: The results of memorization evocation (evoc.) under various dataset-related context, with green and blue curves referring to the memorization difference $\Delta F_{m}$ in the original and TrinEval formats, respectively. In the x-axis, ‘clean’, ‘meta’, ‘dev-fsp’, and ‘seq-fsp’ stand for without dataset-related context, with the name of the dataset, with few-shot prompt from the development set, and with few-shot prompt from the test set ahead of the testing question. These results of $\Delta F_{m}$ indicate the growing memorization effect given the increasing dataset-related information in general. However, the $\Delta F_{m}$ by TrinEval under the strongest memory evocation context remains consistently lower than the one in the original format, e.g., the red dashed line.
In this subsection, we investigate whether the proposed TrinEval reformulation can suppress unnecessary memorization in LLMs, thereby isolating and revealing the models’ genuine capabilities under various circumstances. Following the methodology of Bordt et al. (2024), we introduce dataset-specific cues into the questioning prompts and assess to what extent TrinEval mitigates memorization elicited by such prompts.
To evoke memorization, we embed dataset-related metadata into the input prompts, i.e., the dataset name and in-context few-shot examples drawn from the same dataset (see Appendix E). For this set of experiments, we focus on LLaMA, Mistral, and Vicuna, as their open-source implementations allow access to token-level output probabilities, enabling us to compute the memorization metric $F_{m}$. Since $F_{m}$ lacks a defined absolute zero point indicating complete absence of memorization, we use the $F_{m}$ value obtained from the original MCQ format without any additional context (i.e.,, the vanilla MCQ) as a reference baseline. We then visualize the average change in $F_{m}$ for each memorization-evoking method relative to this baseline.
Specifically, we hypothesize that unintentional data contamination may occur when LLMs are pretrained on datasets scraped from public sources (e.g.,, Hugging Face), where nearby examples from the same dataset may be concatenated during corpus construction. Following the pretraining mechanism of next-token prediction and prior findings from Carlini et al. (2022a), we posit that the inclusion of dataset-similar samples in the prompt may inadvertently trigger memorization. Accordingly, we design a sequence of memorization-evoking perturbations by progressively increasing the contextual cues: (1) providing only the dataset name, (2) adding few-shot examples from the same dataset (i.e.,, samples in the development set or ahead examples in the test set of the same subject of MMLU), and (3) combining both types of context. For each MCQ, we compute the change in $F_{m}$ relative to the reference baseline, capturing the degree to which memorization is elicited by specific context in original or TrinEval formats. Results are illustrated in Fig. 4.
Consistent with the observations by Bordt et al. (2024), the results show that $F_{m}$ increases as more dataset-specific context is introduced, indicating stronger memorization effects. However, across all three tested LLMs, the $\Delta F_{m}$ curve for TrinEval remains consistently lower than that of the original MCQ format (i.e., the red dashed line) regardless of various context used. Notably, even under the strongest memorization-evoking setting, TrinEval’s absolute $F_{m}$ still stays below the vanilla baseline. These findings provide compelling evidence that TrinEval significantly mitigates the influence of memorization, effectively disentangling spurious recall from genuine model capability.
4.4 Q3. TrinEval’s Findings on LLM Memorization and Genuine Capability
(a) $F_{m}$ v.s. $F_{c}$ with Llama on Qwen-extracted triplets.
(b) $F_{m}$ v.s. $F_{c}$ with Mistral on Qwen-extracted triplets.
(c) $F_{m}$ v.s. $F_{c}$ with Vicuna on Qwen-extracted triplets.
(d) $F_{m}$ v.s. $F_{c}$ with Llama on GPT-extracted triplets.
(e) $F_{m}$ v.s. $F_{c}$ with Mistral on GPT-extracted triplets.
(f) $F_{m}$ v.s. $F_{c}$ with Vicuna on GPT-extracted triplets.
@@
Figure 5: The distribution of MCQs based on memorization metric $F_{m}$ vs. capability metric $F_{c}$. According to the values of $F_{m}$ and $F_{c}$, we separate the MCQs into 25 groups and visualize the MCQ distribution from weak to strong with heatmaps.
In this subsection, we aim to explicitly examine the relationship between memorization and genuine problem-solving capability in LLMs, using the metrics $F_{m}$ (memorization) and $F_{c}$ (capability). Since commercial API-based models do not expose full vocabulary-level output probabilities, we focus on open-source LLMs for computing these metrics. After computing the $F_{m}$ and $F_{c}$ scores for all qualified MCQs in TrinEval format, we divide the values of $F_{m}$ and $F_{c}$ into five equal intervals, respectively, resulting in 25 distinct groups of MCQ samples based on their joint distribution. We then use heatmaps to visualize the distribution of samples across these groups, thereby revealing the relationship between memorization and capability for each tested model.
As shown in Fig.5, the majority of MCQs cluster in the lower-left and upper-right corners of the heatmap. This pattern is evident and consistent across our tests. For instance, in the results of LLaMA using Qwen-extracted triplets, 38.57% of MCQs fall within the combined lower-left and upper-right $2\times 2$ grid regions, yielding a Pearson correlation of -0.7755 (p-value $<0.05$) for $F_{m}$ vs. $F_{c}$ of examples within the two square regions; expanding to the $3\times 3$ regions increases coverage to 74.17%, with a stronger correlation of -0.8124 (p-value $<0.05$). Similarly, for Mistral on Qwen triplets, 44.90% of MCQs lie in the two $2\times 2$ corner regions with a correlation of -0.8722 (p-value $<0.05$), and 80.82% in the two $3\times 3$ corner regions with a correlation of -0.8794 (p-value $<0.05$). We take the MCQs within the upper-right 2 X 2 squares in Fig. 5 as the ones memorized by rote and compute the ratio of 19.6% by averaging the percentage of all evaluated LLMs. Similarly, 20.5% is computed as the ratio of MCQs within the lower-left 2 X 3 squares with genuine capability. Additional statistics results are reported in Table 3 in Appendix. These findings consistently indicate a negative correlation between memorization and capability: MCQs with lower memorization scores are more likely to reflect genuine problem-solving capability, whereas high memorization levels correspond to degraded performance, suggesting overfitting or shallow recall.
@@
To interpret these results, we draw an analogy between LLM behavior and the human memory system, which comprises two key components: Short-Term Memory (STM) and Long-Term Memory (LTM) (Shiffrin, 2003). Neurobiological studies reveal that STM relies on transient synaptic protein synthesis with limited temporal persistence and functional scalability. In contrast, LTM is constructed through stabilized neuronal memory traces that constitute an enduring knowledge framework. This neural architecture not only supports STM operations as a cognitive substrate but also enables sophisticated information generalization across diverse contexts. This dichotomy aligns with recent observations in LLMs. As shown by Allen-Zhu and Li (2023) and Ovadia et al. (2023), models trained on corpora with diverse rephrasings exhibit stronger generalization than those trained solely on original formulations. When trained on a single fixed corpus format, LLMs, due to their strong memorization capacity, tend to activate their STM system to memorize at the token level, capturing surface patterns rather than abstract knowledge. In this sense, LLMs are potential rote learners. Reformulations like TrinEval appear to facilitate the encoding of knowledge in a more principled, structured and reusable form, akin to LTM, thereby supporting more generalized and robust evaluation. Further experimental details and results are provided in Appendix C.
Figure 6: Averaged distance of each MCQs between the closest $1\%$ MCQs’ embeddings. ‘Rote Memorization’ refers to MCQ within lower left $2\times 2$ squares that typically exhibits high memorization metric $F_{m}$ and low capability $F_{c}$ while the ‘Genuine Capability Learning’ stands for MCQ lies within the upper right $2\times 2$ squares with low $F_{m}$ but high $F_{c}$. Further results are shown in Appendix F.
To further validate our hypothesis, we analyze the semantic proximity of MCQs within the qualified MMLU dataset by computing their embeddings and measuring the average distance to their closest 1% of samples. Our underlying assumption is that LLMs could better master a knowledge point if it is expressed in various formats in the training corpus. As a result, these knowledge ‘rephrases’ are also encoded densely via LLMs. By selecting only the top 1% nearest neighbors, we aim to retain the samples that are semantically-correlated in knowledge description. We then compare the mean embedding distances of MCQs located in the lower-left (strong memorization, weak capability) and upper-right (weak memorization, strong capability) $2\times 2$ regions of the heatmap in Fig. 5. The corresponding results are presented in Fig. 6. Interestingly, we observe that the average embedding distance among the Genuine Capability Learning MCQs is much lower than (nearly half) that of the Rote Memorization MCQs. This suggests that memorized MCQs are more sparsely distributed in the embedding space, while non-memorized, capability-driven MCQs tend to form tighter semantic clusters. This observation also aligns well with the cognitive analogy of STM and LTM: rote memorization leads to fragmented, context-specific encoding, whereas genuine capability emerges from structured, reusable representations. The results of alternative thresholds are shown in Table 5. Though it is well believed that memorization may lead to better but cheating performance of LLMs, we prove that the more LLMs memorize, the worse they are at solving problems.